Please note: You are viewing the unstyled version of this website. Either your browser does not support CSS (cascading style sheets) or it has been disabled. Skip navigation.

June 9, 2014 - ESP Update

   Print Change text to small (default) Change text to medium Change text to large

Here's the latest ESP results as of Monday, June 9, 2014...

All 3 ESPs worked today and reported positive hits for A. fundyense as we start the 6th week of operations.

ESP-1 (don) near NERACOOS B reported almost 400 cells/L today (June 9), down from the 500-700 cells/L range reported last week.

ESP-2 (jake) in Casco Bay was in action today and reported a solid signal in the range of 400-500 cells/L, back up from a dip last week.

ESP-3 (dennis) off of Pemaquid Point also reported a positive signal in the 100-200 cells/L range after being totally blanked last Friday.

With a final ground truth survey scheduled starting June 15, we're hoping the batteries hold on...power usage estimates indicate we have another 10 days or so of operations remaining.

Will send the latest update each day this week,

Bruce Keafer



Last updated: July 1, 2019
 


whoi logo

Copyright ©2007 Woods Hole Oceanographic Institution, All Rights Reserved, Privacy Policy.
Problems or questions about the site, please contact webdev@whoi.edu